# Detection and Characterization of Coordinated Online Behavior: A Survey

### v1

@@ Abstract.
Coordinated behavior, coordination, synchronization, online social networks, network science
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
With the advent of social media platforms, coordination has also become a fundamental component of online interactions. Social media users are now provided with a broad array of tools to coordinate with each other, such as hashtags that enable them to collectively discuss specific topics [122, 13]. Online platforms have become a suitable environment for organizing social and political movements worldwide, giving rise to phenomena such as online activism [90, 99], boycotts [10, 71], and protests [75, 119]. The 2011 Arab Springs are a notable example, being largely organized through social media [55]. At the same time however, scholars found evidence of online coordination being exploited by nefarious actors for all sorts of malicious purposes. For instance, disinformation campaigns often leverage actors that coordinate their actions to maximize the outreach of their false narratives [118, 114, 63, 131]. Similarly, coordination is employed within information operations [19, 131, 96] and astroturfing, which involves creating the false appearance of grassroots support for a target cause, product, or person [63]. Also social bots and trolls exploit online coordination to amplify messages, manipulate trends, or spread disinformation [86, 83, 72, 142]. Finally, coordination also results from, and contributes to, the formation of echo chambers and online polarization [132]. Figure 2 highlights the complex, multifaceted, and intertwined nature of online coordination, which represents the underlying substrate of many diverse yet interconnected phenomena that permeate online social environments.
@@
Figure 2. Number of articles published yearly on coordinated online behavior. A steep rise is observed after Facebook introduced CIB in 2018.
Recognizing the profound impact of online coordination on social media as well as its consequences on the offline world, both researchers [131, 101, 139, 99] and industry stakeholders [36, 40, 37] devoted a great deal of efforts to study its dynamics and to develop effective strategies to detect and mitigate its malicious instances. Research sped up significantly after 2018, when Facebook introduced the concept of coordinated inauthentic behavior (CIB) [36], marking a milestone in the development of the field. As shown in Figure 2, subsequent years saw a surge of interest on the subject, testified by the steadily growing number of published papers. This survey is motivated by this thriving interest on online coordination, which resulted in the availability of a large body of work. However, while Facebook’s interest towards online coordination was constrained to inauthentic behaviors as a response to the threat of orchestrated campaigns [36], here we embrace a more holistic and unbiased view by focusing on the broader concept of coordinated behavior. This inclusive approach allows for the analysis of a broader spectrum of works, including those focused on legitimate collective actions, offering a more comprehensive understanding of the coordination dynamics that shape digital spaces and fostering nuanced perspectives that go beyond mere threat detection. In spite of the many efforts, the existing literature still reflects the complexities and ambiguities surrounding this phenomenon. Among them is the limited agreement on a shared definition, which also hinders operationalization. Complexity also emerges from the diversity of methods proposed for detecting and characterizing coordinated online behavior, which impairs comparisons between different works and limits the generalizability of findings. Finally, the use of the same coordination technique by actors with disparate motivations poses challenges to estimating the impact and effects of online coordination.
@@
Comprehensively modeling coordinated online behavior has far-reaching implications. On a theoretical level, it reconciles diverse definitions and provides a foundational framework for future research. On a technical level, it critically evaluates and categorizes existing detection and characterization methods, informing the development of the next generation of robust and adaptive tools for studying both malicious and neutral instances of coordinated behavior. By shedding light on online coordination—a fundamental dynamic of computer-mediated human behavior—this survey contributes to safeguarding online integrity and to fostering positive interactions in digital spaces. It offers valuable insights for shaping future methodologies, platforms, and policies, and it also contributes to enriching the interdisciplinary research occurring at the intersection between computer science and social dynamics.
Scope.
Coordinated online behavior is orthogonal to many of the research topics tackled in fields such as Web and social media analysis, online social networks security, as well as social computing and computational social science, as shown in Figure 2. Therefore, a large number of works from these scientific communities implicitly or explicitly deal with online coordination. However, this survey is constrained to those papers that address online coordination explicitly and that provide relevant contributions to its detection, characterization, or understanding. In practice, we identified an initial set of candidate papers by selecting from Google Scholar and Scopus all those papers including the term “behavior” and at least one term among “coordinated”, “inauthentic”, “collaborative”. Each candidate paper was then evaluated by one of the authors of this survey to filter out unrelated works. Furthermore, we also manually checked all references from each of the related works retained after the previous filtering step, in order to identify possible additional related works. The final set of works categorized and critically discussed in this survey includes 84 papers published between 2014 and 2024, as shown in Figure 2.
Organization.
@@
This survey is structured as follows. Section 2 presents the theoretical foundations of coordinated online behavior, proposing a new general definition and laying out a comprehensive conceptual framework. Section 3 bridges the theoretical and methodological parts by defining the detection and characterization tasks. Section 4 discusses the existing literature on the detection task, while Section 5 focuses on the characterization task. Section 6 summarizes the main outstanding challenges and suggests promising directions of research. Finally, Section 7 concludes the survey.
@@ 2.1. Offline coordination
Definition 2.1 denotes coordination as the organizational overhead that multiple actors incur into when pursuing goals together. We note that this and similar definitions [77, 80, 7, 6] implicitly leverage the fundamental components of coordination, which we explicitly define as follows:
Definition 2.1. Coordination components: A set of two or more actors who perform activities in order to achieve goals.
Definition 2.1 introduces the fundamental components of coordination: actors, activities, and goals. Being coordination a nuanced concept, the theoretical modeling of these components can have major implications on downstream analyses and results. For example, in the case of communities or groups of users, each user in the group can be treated as a standalone actor, or alternatively the entire group may be considered as a single actor. Similar choices must be made when modeling the activities that allow actors to coordinate. Each actor typically performs multiple activities during any given time frame, and each of these activities might contribute differently to the overall coordination. Therefore, the choice of activities to model during an analysis directly impacts the resulting observed coordination [79]. Finally, the analyst is often interested in knowing the goal that the actors pursue when performing the activities. However, the actors may not all have the same goal, or even have any explicit goal at all [78]. These reflections on the components of coordination surface some of the challenges that early scholars faced since the 80s when studying offline coordination. Interestingly, many of these challenges carry over to the study of online coordination, informing the development of a new comprehensive definition and conceptual framework.
@@ 2.2. Concepts and definitions by online platforms
2.2.1. Facebook/Meta
After the public disclosure that the Internet Research Agency (IRA) had strategically exploited the platform to influence the 2016 US presidential election [100], Facebook began publishing reports detailing how their services were abused and the actions taken in response. In unveiling further actions against the IRA, in July 2018 Facebook introduced the concept of coordinated inauthentic behavior (CIB) [37]. A few months later, they supplied it with a first definition, and in October 2019 with a second one.
@@ 2.2.2. Twitter/X
In October 2018, the platform released a public archive containing data about identified information operations (IOs). [^1] https://transparency.twitter.com/en/reports/moderation-research.html (accessed: 31/07/2024) Although not explicitly stated in Twitter’s definition at the time, a certain degree of coordination is necessary for the success of an IO [131, 19]. However, it was not until January 2021 that Twitter adopted a similar approach to Facebook and released a definition that explicitly addresses instances where coordination is leveraged to cause harm both online and offline.
@@ 2.2.5. Ambiguities and limitations
This brief survey of the main concepts and definitions proposed by social media shows that online platforms are at the forefront in the analysis of coordinated online behavior. However, their conceptualizations are driven primarily by pressing practical regulation needs and by immediate contingencies, rather than by methodological rigor and theoretical soundness [30]. Faced with specific instances of malicious coordination, online platforms adopt different and at times contrasting definitions, adding confusion and ambiguity to the already challenging task of defining an inherently nuanced and complex phenomenon. The current tangled landscape of platform definitions means that certain coordinated efforts may be categorized as such by some platforms but not by others, contingent upon the concepts they adhere to. What may be identified as coordinated behavior on one platform could be overlooked or dismissed on another, leading to inconsistencies in detection and mitigation. As a notable example, in June 2020 many teenagers organized on TikTok to reserve tickets for a Donald Trump rally to be held in Tulsa, OK (US). By mass-reserving and later cancelling their participation, they prevented others from making reservations and artificially inflated expected attendance numbers, ultimately causing an embarrassing number of empty seats at the rally [31]. Facebook’s head of security commented that, while tactical and sophisticated, they would have not acted upon this campaign as an instance of CIB (as per Definition 2.3), since it did not make use of fake accounts nor aimed to mislead Facebook users [31]. A similar case is the Chinese Spamouflage campaign, a cross-platform propaganda effort against the Hong Kong pro-democracy movement [98]. Various platforms reported takedowns of Spamouflage instances, with Google and Twitter labeling it as CIO and IO (as per Definition 2.6), respectively. However, Reddit interpreted Spamouflage differently, recognizing its low-quality and one-sided content, but refraining from considering these activities as rule violations. [^2] https://www.reddit.com/r/redditsecurity/comments/dp9nbg/reddit_security_report_october_30_2019/ (accessed: 31/07/2024) Evidently, platforms address instances of coordinated online behavior disparately, and the criteria to establish the legitimacy of user behaviors lack objectivity. These discrepancies underscore the need for a unified understanding and standardized approach to defining and addressing online coordination, one that transcends platform-specific idiosyncrasies and fosters a more cohesive response to this challenge.
@@ 2.3. Concepts and definitions in academic literature
As shown in Figure 2, scientific interest in coordinated online behavior has drastically risen in the last few years. However, akin to social media platforms, scholars have generally refrained from proposing theoretically-grounded and general definitions. The majority of the existing studies adopt Facebook’s Definition 2.2 of CIB [85, 87]. Instead, those who propose their own conceptualization mainly provide operational definitions useful for the development of coordination detection methods [75, 99, 102]. Table 1 reports some examples of operational definitions proposed recently. As shown, no definition is comprehensive and general enough to adequately describe the multifaceted phenomenon of coordinated online behavior. The existing conceptualizations of the phenomenon appear to be influenced by the specific technique employed for its detection. Consequently, the criteria used to define coordination vary, encompassing aspects ranging from similar behavior [99, 20, 74, 138, 75, 102] to synchronicity [74, 75, 143] and common intent [56, 63]. As practical examples of the above, some definitions and the resulting detection methods revolve around the anomalous use of hashtags or URLs by multiple users [48, 75], or repeated screen name swapping [102]. However, while no single definition comprehensively describes online coordination, each grasps one or more relevant properties, as highlighted in table. In turn, these properties can inform a general definition coordinated online behavior.
@@ 2.4. General definition and fundamental components of coordinated online behavior
Definition 2.2. Coordinated online behavior: $\underbrace{\text{A group of users}}_{\textrm{{actors}}}$ $\underbrace{\text{who perform synergic actions}}_{\textrm{{actions}}}$ $\underbrace{\text{in pursuit of an intent}}_{\textrm{{intent}}}$.
Definition 2.4 delineates coordinated online behavior based on three fundamental components—actors, actions, and intent—that are similar to the components of offline coordination outlined in Definition 2.1. Definition 2.4 and its three fundamental components enable the comprehensive mapping of all instances of online coordination, as discussed in the following.

### v2

@@ Abstract.
copyright: none
@@ 1. Introduction
With the advent of social media platforms, coordination has also become a fundamental component of online interactions. Social media users are now provided with a broad array of tools to coordinate with each other, such as hashtags that enable them to collectively discuss specific topics [129]. Online platforms have become a suitable environment for organizing social and political movements, giving rise to phenomena such as online activism [93, 100], boycotts [68], protests [73, 126], and coordinated crisis communication by democratic institutions during emergencies [151]. The 2011 Arab Springs are a notable example, being largely organized through social media [126]. At the same time however, scholars found evidence of online coordination being exploited by nefarious actors for all sorts of malicious purposes. For instance, disinformation campaigns often leverage actors that coordinate their actions to maximize the outreach of their false narratives [125, 59, 138]. Similarly, coordination is employed within information operations [18, 138, 99], coordinated social media manipulation [87, 130], and astroturfing, which involves creating the false appearance of grassroots support for a target cause, product, or person [59]. Also, social bots (automated accounts [25, 96]) and trolls (human-driven accounts engaging in collective disruptive or deceptive behavior [153]) exploit coordination to amplify messages, manipulate trends, or spread disinformation [82, 69, 153]. Coordination both results from and contributes to echo chambers and online polarization [139]. Figure 2 illustrates its multifaceted nature, underlying diverse and interconnected phenomena in online social media.
@@
Figure 2. Number of articles published yearly on coordinated online behavior.
Recognizing the profound impact of online coordination on social media as well as its consequences on the offline world, both researchers [138, 102, 146, 100] and industry stakeholders [35] devoted a great deal of efforts to study its dynamics and to develop effective strategies to detect and mitigate its malicious instances. Research sped up significantly after 2018, when Facebook introduced the concept of coordinated inauthentic behavior (CIB) [35], marking a milestone in the development of the field. As shown in Figure 2, subsequent years saw a surge of interest on the subject, testified by the steadily growing number of published papers, with only a minor fluctuation in 2023. This survey is motivated by this thriving interest on online coordination, which resulted in the availability of a large body of work. However, while Facebook’s interest towards online coordination was constrained to inauthentic behaviors as a response to the threat of orchestrated campaigns [35], here we embrace a more holistic and unbiased view by focusing on the broader concept of coordinated behavior. This inclusive approach allows for the analysis of a broader spectrum of works, including those focused on legitimate collective actions, offering a more comprehensive understanding of the coordination dynamics that shape digital spaces and fostering nuanced perspectives that go beyond mere threat detection. In spite of the many efforts, the existing literature still reflects the complexities and ambiguities surrounding this phenomenon. Among them is the limited agreement on a shared definition, which also hinders operationalization. Complexity also emerges from the diversity of methods proposed for detecting and characterizing coordinated online behavior, which impairs comparisons between different works and limits the generalizability of findings. Finally, the use of the same coordination technique by actors with disparate motivations poses challenges to estimating the impact and effects of online coordination.
@@
Significance. Comprehensively modeling coordinated online behavior has far-reaching implications. On a theoretical level, it reconciles diverse definitions and provides a foundational framework for future research. On a technical level, it critically evaluates and categorizes existing detection and characterization methods, informing the development of the next generation of robust and adaptive tools for studying both malicious and neutral instances of coordinated behavior. By shedding light on online coordination—a fundamental dynamic of computer-mediated human behavior—this survey contributes to safeguarding online integrity and to fostering positive interactions in digital spaces. It offers valuable insights for shaping future methodologies, platforms, and policies, and it also contributes to enriching the interdisciplinary research occurring at the intersection between computer science and social dynamics.
Scope. Coordinated online behavior is orthogonal to many of the research topics tackled in fields such as Web and social media analysis, online social networks security, as well as social computing and computational social science, as shown in Figure 2. Therefore, a large number of works from these scientific communities implicitly or explicitly deal with online coordination. While coordinated behavior can manifest across a wide range of online environments (e.g., e-commerce platforms, review systems, or online marketplaces [105]), this survey focuses primarily on online social media platforms, where coordination has been most extensively studied and where interactions are more readily observable at scale. Accordingly, this survey is constrained to those papers that address online coordination explicitly within this context and that provide relevant contributions to its detection, characterization, or understanding.
Identification
Screening
Included
Records identified from:Scopus ($n=353$)Google Scholar ($n=1000$) Additional studies identified throughbackward reference searching($n=38$) Records screened ($n=1206$) Studies included in review ($n=122$) Records removed before screening:Duplicate records removed ($n=147$) Full-text articles excluded ($n=1123$):Not in English ($n=4$)Out of scope ($n=1118$) Identification of studies via databases and registers Identification of studies via other methods
Figure 3. PRISMA diagram summarizing the stages used to construct the final corpus of studies considered in this survey.
Figure 4. Distribution of papers in the dataset according to the disciplinary area of the publication venue, highlighting the predominance of Computer Science venues alongside contributions from Social Sciences and interdisciplinary outlets.
Literature search and selection. A systematic literature search was conducted to identify studies addressing coordinated behavior on social media platforms using two databases: Scopus and Google Scholar. The Scopus query targeted papers containing the following search strings:
(“coordinated behavior” OR “coordinated behaviour” OR “coordinated activity” OR “coordinated campaign” OR “coordinated inauthentic behavior” OR “coordinated inauthentic behaviour” OR “inauthentic behavior” OR “inauthentic behaviour” OR “information operation*”) AND (“social media” OR online OR twitter OR facebook OR instagram OR reddit OR telegram OR tiktok OR youtube)
The search was restricted to publications between 2014 and 2026, written in English, and within relevant subject areas (Computer Science, Social Sciences, Decision Sciences, Engineering, and Multidisciplinary). This query returned 353 records. To broaden coverage, a complementary search was performed on Google Scholar using similar search strings, collecting the first 1000 results returned by the platform. This choice aims to maximize coverage while reducing potential bias introduced by the platform’s ranking mechanism. After merging the two sources, 147 duplicate records were removed, leaving 1206 records for screening. Titles and abstracts were manually reviewed, leading to the exclusion of 4 non-English papers and 1118 records that were out of scope (e.g., biological coordination, general sociological coordination, or studies on bot detection or information operations not explicitly addressing coordinated behavior). Most of these discarded records derived from the search on Google Scholar, likely due to the choice of selecting the first 1000 results to maximize the coverage. After the screening phase, 83 papers were retained. A backward reference search of these studies identified an additional 38 relevant papers, resulting in a final corpus of 122 papers included in this survey. The full selection workflow is illustrated in the PRISMA diagram in Figure 4.
Subject area distribution. To analyze the disciplinary distribution of the collected studies, each paper was assigned a subject area based on the official classification associated with the publication outlet. For journal articles, the subject area was determined using the Scimago Journal Rank classification system [^1] https://www.scimagojr.com. For conference papers, the classification provided by the CORE conference ranking portal was used [^2] https://portal.core.edu.au/conf-ranks/. For preprints available on arXiv, the category assigned by arXiv itself was adopted. The resulting distribution of papers across subject areas is shown in Figure 4.
@@
Organization. This survey is structured as follows. Section 2 presents the theoretical foundations of coordinated online behavior, proposing a new general definition and laying out a comprehensive conceptual framework. Section 3 bridges the theoretical and methodological parts by defining the detection and characterization tasks. Section 4 discusses the existing literature on the detection task, while Section 5 focuses on the characterization task. Section 6 summarizes the main outstanding challenges and suggests promising directions of research. Finally, Section 7 concludes the survey.
@@ 2.1. Offline coordination
Definition 2.1 denotes coordination as the organizational overhead that multiple actors incur into when pursuing goals together. We note that this and similar definitions [77, 76, 9, 10] implicitly leverage the fundamental components of coordination, which we explicitly define as follows: Definition 2.1. Coordination components: A set of two or more actors who perform activities in order to achieve goals. Definition 2.1 introduces the fundamental components of coordination: actors, activities, and goals. Being coordination a nuanced concept, the theoretical modeling of these components can have major implications on downstream analyses and results. For example, in the case of communities or groups of users, each user in the group can be treated as a standalone actor, or alternatively the entire group may be considered as a single actor. Similar choices must be made when modeling the activities that allow actors to coordinate. Each actor typically performs multiple activities during any given time frame, and each of these activities might contribute differently to the overall coordination. Therefore, the choice of activities to model during an analysis directly impacts the resulting observed coordination [75]. Finally, the analyst is often interested in knowing the goal that the actors pursue when performing the activities. However, the actors may not all have the same goal, or even have any explicit goal at all [78]. These reflections on the components of coordination surface some of the challenges that early scholars faced since the 80s when studying offline coordination. Interestingly, many of these challenges carry over to the study of online coordination, informing the development of a new comprehensive definition and conceptual framework.
@@ 2.2. Concepts and definitions by online platforms
2.2.1. Meta (Facebook, Instagram, Threads)
After the public disclosure that the Internet Research Agency (IRA) had strategically exploited the platform to influence the 2016 US presidential election [155, 146, 118], Facebook began publishing reports detailing how their services were abused and the actions taken in response. In unveiling further actions against the IRA, in July 2018 Facebook introduced the concept of coordinated inauthentic behavior (CIB) [35]. A few months later, they supplied it with a first definition, and in October 2019 with a second one. These definitions, originally introduced by Facebook, also applied to Instagram due to their shared ownership. Then they have been extended across Meta’s platforms, including Facebook, Instagram, and Threads.
@@ 2.2.2. Twitter/X
In October 2018, the platform released a public archive containing data about identified information operations (IOs) [117]. Although not explicitly stated in Twitter’s definition at the time, a certain degree of coordination is necessary for the success of an IO [138, 18]. However, it was not until January 2021 that Twitter adopted a similar approach to Facebook and released a definition that explicitly addresses instances where coordination is leveraged to cause harm both online and offline.
@@ 2.2.4. Reddit
2.2.5. Tiktok
In 2023, TikTok revisited Facebook’s definition of CIB, explicitly linking coordination with inauthenticity. In doing so, it emphasizes coordinated activity aimed at misleading users or influencing public discourse. Among the cited examples, TikTok highlights coordination in the context of information operations, such as accounts covertly promoting political candidates or issues, or posting content on behalf of foreign entities without proper disclosure.
Definition 2.9 (Covert Influence Operations).
2023 Coordinated, inauthentic behaviors where networks of accounts work together to mislead people or our systems and try to strategically influence public discussion. This may include attempting to undermine the results of an election, influencing parts of an armed conflict, or shaping public discussion of social issues. [131].
@@ 2.2.6. Ambiguities and limitations
This brief survey of the main concepts and definitions proposed by social media shows that online platforms are at the forefront in the analysis of coordinated online behavior. However, their conceptualizations are driven primarily by pressing practical regulation needs and by immediate contingencies, rather than by methodological rigor and theoretical soundness [35]. Faced with specific instances of malicious coordination, online platforms adopt different and at times contrasting definitions, adding confusion and ambiguity to the already challenging task of defining an inherently nuanced and complex phenomenon [89].
The current tangled landscape of platform definitions means that certain coordinated efforts may be categorized as such by some platforms but not by others, contingent upon the concepts they adhere to. What may be identified as coordinated behavior on one platform could be overlooked or dismissed on another, leading to inconsistencies in detection and mitigation. As a notable example, in June 2020 many teenagers organized on TikTok to reserve tickets for a Donald Trump rally to be held in Tulsa, OK (US). By mass-reserving and later cancelling their participation, they prevented others from making reservations and artificially inflated expected attendance numbers, ultimately causing an embarrassing number of empty seats at the rally [31]. Facebook’s head of security commented that, while tactical and sophisticated, they would have not acted upon this campaign as an instance of CIB (as per Definition 2.3), since it did not make use of fake accounts nor aimed to mislead Facebook users [31]. A similar case is the Chinese Spamouflage campaign, a cross-platform propaganda effort against the Hong Kong pro-democracy movement. [^3] https://graphika.com/reports/spamouflage-breakout (accessed: 31/07/2024) Various platforms reported takedowns of Spamouflage instances, with Google and Twitter labeling it as CIO and IO (as per Definition 2.6), respectively. However, Reddit interpreted Spamouflage differently, recognizing its low-quality and one-sided content, but refraining from considering these activities as rule violations. [^4] https://www.reddit.com/r/redditsecurity/comments/dp9nbg/reddit_security_report_october_30_2019/ (accessed: 31/07/2024)
Evidently, platforms address instances of coordinated online behavior disparately, and the criteria to establish the legitimacy of user behaviors lack objectivity. These discrepancies underscore the need for a unified understanding and standardized approach to defining and addressing online coordination, one that transcends platform-specific idiosyncrasies and fosters a more cohesive response to this challenge.
@@ 2.3. Concepts and definitions in academic literature
As shown in Figure 2, scientific interest in coordinated online behavior has drastically risen in the last few years. However, akin to social media platforms, scholars have generally refrained from proposing theoretically-grounded and general definitions. The majority of the existing studies adopt Facebook’s Definition 2.2 of CIB [85, 86]. Instead, those who propose their own conceptualization mainly provide operational definitions useful for the development of coordination detection methods [73, 100, 103]. Table 1 reports some examples of operational definitions proposed recently. As shown, no definition is comprehensive and general enough to adequately describe the multifaceted phenomenon of coordinated online behavior. The existing conceptualizations of the phenomenon appear to be influenced by the specific technique employed for its detection. Consequently, the criteria used to define coordination vary, encompassing aspects ranging from similar behavior [100, 19, 72, 145, 73, 103] to synchronicity [72, 73, 155] and common intent [51, 59]. As practical examples of the above, some definitions and the resulting detection methods revolve around the anomalous use of hashtags or URLs by multiple users [44, 73], or repeated screen name swapping [103].
More fundamentally, the field still lacks a formal and statistically grounded definition of coordination, including well-defined null models against which coordinated behavior can be rigorously assessed. As a consequence, many approaches rely on platform- and task-specific heuristics, which complicates cross-domain comparisons, reproducibility, and the estimation of effect sizes. This limitation echoes challenges observed in related areas such as bot detection, where simplistic data collection and labeling practices have been shown to hinder generalization even across datasets from the same platform [50]. While no single definition fully captures the complexity of coordinated online behavior, existing works collectively highlight key properties of the phenomenon. These properties can serve as building blocks toward a more general and theoretically grounded definition of coordination.
@@ 2.4. General definition and fundamental components of coordinated online behavior
Definition 2.2. Coordinated online behavior: $\underbrace{\text{A group of users}}_{\textrm{{actors}}}$ who perform synergic $actions\underbrace$$\text${who perform synergic $_{\textrm{{actions}}}$ in pursuit of an $intent\underbrace$$\text${in pursuit of an $_{\textrm{{intent}}}$. Definition 2.4 delineates coordinated online behavior based on three fundamental components—actors, actions, and intent—that are similar to the components of offline coordination outlined in Definition 2.1. Definition 2.4 and its three fundamental components enable the comprehensive mapping of all instances of online coordination, as discussed in the following.